Circular No. 5818

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1993U AND 1993V IN ANONYMOUS GALAXIES C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of two

apparent supernovae found on a single B plate taken June 16.9 UT with the OCA Schmidt telescope by D. Albanese and himself. The supernova candidates also appear on a film taken on June 19.9. SN 1993U (magnitude B_{j} ∼ 19.5) is located at α = 15^{h}23^{m}44^{s}.62, δ = +50^{o}45′07′′.4 (equinox 1950.0), which is 8′′.4 east and 6′′.6 north of the host galaxy’s nucleus, the galaxy having m_{pg} ∼ 18.5. There is a nearby star of m_{pg} ∼ 14 with end figures 44^{s}.81, 55′′.5. SN 1993V (B_{j} ∼ 19) is located at α = 15^{h}29^{m}32^{s}.06, δ = +50^{o}27′56′′.6, which is 6′′.9 west and 16′′.3 south of the nucleus of a galaxy of m_{pg} ∼ 18. A nearby star of B ∼ 16 has end figures 27^{s}.44, 26′56′′.2. Palomar Sky Survey prints show no objects to blue mag 20.5 and to red mag 20 for both supernovae. M. Dennefeld obtained several CCD images on June 21 with the 1.2-m telescope at Observatoire de Haute-Provence under nonphotometric conditions, showing SN 1993U to be very blue and SN 1993V to be red.

SUPERNOVA 1993R IN NGC 7742 B. Leibundgut and W. Vacca, University of California Berkeley, report

that preliminary inspection of unreduced spectrograms (range 350-510 and 540-1000 nm), obtained on June 19.45 UT with the 3-m Shane Telescope at Lick Observatory, show SN 1993R (cf. IAUC 5812) in an advanced stage. Broad nebular lines of Ca II (850 nm), [Ca II] (730 nm) and [O II] (732 nm), [O I] (630 nm), and [O III] (501 nm) are seen as strong emissions. Broad Hα is also detected, indicating that this supernova is of type II. The line is, however, relatively weak compared to the other emissions and blended with narrow emission from the galaxy.

NOVA AQUILAE 1993 Further photoelectric photometry (comparison star HR 7313, cf. IAUC

5794): May 26.285 UT, V = 8.00, B–V = +0.44 (D. H. Kaiser, Columbus, IN); 27.267, 8.11, +0.41 (Kaiser); June 18.99, 10.16 ± 0.03, +0.22 ± 0.05 (H. Mikuž, Ljubljana, Slovenia); 19.94, 10.29 ± 0.03, – (Mikuž). Kaiser notes that SAO 124400, the 9th-magnitude star to the nova’s west, varies by at least 0.05 mag over 9 days, though it has evidently not been reported previously as variable.

1993 June 21 Daniel W. E. Green